# Exact Infrared Triangle in Massless sQED with Long-range Interactions

### v1

@@ 1 Introduction
The simplest example of such an infrared triangle consists of the soft photon theorem in QED, the electromagnetic kick memory effect and large gauge symmetry which acts on matter as a local enhancement of global phase rotations, or, superphaserotations; it manifests itself in the conservation of a Noether charge at every angle whose global version is the familiar U(1) charge of QED [He:2014cra, Lysov:2014csa]. Analogous infrared triangles have been found in other theories including gravity [Strominger:2013jfa], non-Abelian gauge theory [He:2015zea] and supersymmetric theories [Strominger:2013lka]. The study of the infrared structure has also been extended to subleading orders in the low-energy expansion [Lysov:2014csa, Campiglia:2015kxa, Kapec:2014opa, Campiglia:2016efb] which translates to weaker large-distance fall-offs of the radiative field in the corresponding memory effect analyses and to overleading orders in a large-distance expansion of the associated symmetry parameters [Campiglia:2016hvg, Campiglia:2016efb, Choi:2024ygx, Choi:2024ajz].
The long-range nature of electromagnetic and gravitational interactions turns out to modify all these infrared triangles beyond the leading one: early/late-time divergences were shown to dress the matter fields [Campiglia:2019wxe] by logarithmically divergent phases and this behavior dominates over the free field asymptotics! Only recently have these effects been accounted for in the discussion of the infrared triangles [Choi:2024ygx, Choi:2024mac, Choi:2024ajz] and so far only in (scalar) QED and gravity with massive matter. Our goal in this work is to connect the logarithmic soft photon theorem in massless scalar QED to an asymptotic symmetry analysis. A corresponding analysis for gravity coupled to massless matter is in progress [Choi-Kadhe-Puhm].
@@
Unlike the leading soft factor (1.2), the subleading soft factor $S_{0}$ is not universal. It receives corrections from a (finite) set of effective field theory operators [Elvang:2016qvq] with three-point interactions as well as loop corrections [Bern:2014oka, Laddha:2018myi, Laddha:2018vbn, Sahoo:2018lxl]. These loop corrections enter at $O(\ln\omega)$ and so are overleading compared to the subleading tree-level soft factor $\sim\omega^{0}$ which the presence of the logarithm renders ambiguous.
@@
In (massive) scalar QED, it has been conjectured [Karan:2025ndk] that given a well-defined soft expansion to all orders in the loop expansion, the ratio between the $(N+1)$-particle amplitude including a soft photon and the corresponding $N$-particle amplitude without that soft photon is infrared-finite. This allows us to define the loop-corrected expansion of the soft factor as
@@
The logarithmic soft theorem in theories with massive matter receives contributions from classical infrared and quantum loop effects [Sahoo:2018lxl]
@@
An asymptotic symmetry explanation of the classical soft theorem was given in [Choi:2024mac] which derived a ‘soft’ charge that maps on to the soft photon insertion in the S-matrix and a ‘hard’ charge whose action on the matter fields produces the classical soft factor. The conservation of the sum of these logarithmic soft and hard charges between past and future boundaries is exact in the electromagnetic coupling and yields the classical logarithmic soft photon theorem which is one-loop exact.
Here we would like to explain the symmetry underlying the classical logarithmic soft photon theorem in theories with massless matter. However, it is easy to see that the (naive [^2] The absence of potential collinear divergences is shown in appendix D.) massless limit of (1.7) vanishes. This points towards the classical one-loop corrected scattering structure of massless QED being trivial: long-range effects do not correct the subleading tree-level structure. Moreover the subleading tree-level soft theorem is not rendered ambiguous as they would have been in the presence of logarithms. The only non-vanishing contribution comes from quantum effects, but how to explain them from an asymptotic symmetry analysis is as of yet unclear, and we will not attempt to do so here [^3] Attempts in this direction were made in [Campiglia:2019wxe].. In an asymptotic symmetry derivation of the classical logarithmic soft theorem with massless matter we thus have to show that both soft and hard charges exactly vanish to all orders in the electromagnetic coupling!
@@
As ansatz for the relevant symmetry transformation we consider the same linearly divergent superphaserotation that is responsible for the subleading tree-level soft photon theorem. This is motivated by the corresponding derivation of the logarithmic soft photon theorem in massive scalar QED. Indeed we will see here that this ansatz reproduces the leading and subleading tree-level soft theorems in massless scalar QED in addition to explaining the exact vanishing of logarithmic soft photon theorem from symmetry [^4] See also the related work [AtulBhatkar:2019vcb] .. Moreover, we determine the long-range correction to the charge associated to the subleading soft theorem [Campiglia:2016hvg], at higher order in the electromagnetic coupling which corresponds to the one-loop correction to Low’s subleading soft photon theorem. Care has to be taken with potential collinear divergences, which we show do, however, not affect our classical analysis.
@@ 2.2 Interacting fields
Because the correction to the Klein-Gordon equation is subleading compared to the free equation, the interacting scalar field is still dominated by the free behaviour, $\phi(x)=\phi^{\rm free}(x)+o(1/r)$, which implies $\phi_{0,n}=\delta_{n,0}\phi_{0,0}$ in (2.26). This behavior is in stark contrast to the situation in massive QED [Choi:2024mac] where Coulombic modes give rise to a logarithmically divergent dressing for the massive scalar field. In massless QED, the matter dressing by Coulombic gauge modes only comes in at subleading order
@@ 3 Superphaserotation symmetry
are referred to as large if they give rise to non-vanishing Noether charges on the asymptotic phase space of the theory. To avoid any potential confusion with gauge redundancies, we follow [Choi:2024mac] and refer to these physical transformations by how they act on matter fields, namely as a local enhancement of global phase rotations, or, superphaserotations.
@@
The solution to this equation that is relevant for our discussion is the one that diverges linearly in $r$ [Campiglia:2016hvg],
@@
where ${\overline{\epsilon}(\hat{x})}$ is a free function on $S^{2}$. Asymptotic charges with such divergent symmetry parameters have been found in [Campiglia:2016hvg] to account for Low’s subleading tree-level soft photon theorem [Low:1958sn]. Accounting for the long-range interactions, it has been shown [Choi:2024mac] that the same symmetry transformation governs the logarithmic soft photon theorem in massive scalar QED. It is thus natural to assume that the same symmetry that governs the massless limit of Low’s soft theorem is also responsible for the logarithmic soft photon theorem in massless scalar QED.
@@ 4 Superphaserotation charge
To make the connection with soft theorems, it is useful to identify the ‘soft’ and ‘hard’ contributions to this charge (4.8), i.e. the contributions that will map onto the soft photon insertion in the S-matrix and the contributions coming from the superphaserotation of the matter fields which will map onto the soft factor. Interestingly, unlike in massive scalar QED, the term $\text{d}\epsilon\wedge\star{F}$ in (4.8) turns out to contribute not just to the soft charge but also to the hard charge. This is already the case for the charge constructed in [Campiglia:2016hvg], whose conservation was shown to be equivalent to the subleading tree-level soft photon theorem in massless scalar QED. Here we will find a similar result for the logarithmic soft photon theorem.
@@
As in the massive QED analysis [Choi:2024mac] it turns out the tower of logarithmic terms in (4.11) vanishes. Thus, the charge (4.10), computed from the symplectic structure is linearly divergent in $r$ and $u$, logarithmically divergent in $r$, and potentially logarithmically divergent in $|u|$, as arising from the $1/|u|$ term in (4.11). To regulate the logarithmic divergences at large $r$ and large-$|u|$ we introduce a cutoff [^5] Note that choosing different cutoffs to regulate large-$r$ and large-$u$ divergences only affects the charge corresponding to the subleading tree-level soft theorem. This ambiguity is expected: logarithmic terms in the soft expansion render the power-law tree-level expansion (beyond the Weinberg pole) ambiguous. at $\ln\Lambda^{-1}$, similar to the case of massive QED [Choi:2024mac]. In the following we will show that the three charges, $Q^{(-1)}$, $Q^{(0)}$ and $Q^{(\ln)}$, that we can extract from (4.10) are associated to, respectively, the leading Weinberg, subleading Low, and logarithmic Sahoo-Sen soft theorems.
@@ 4.3 Loop-corrected subleading charge $Q^{(0)}$
The covariant form of this expression matches the one obtained in [Campiglia:2016hvg] whose conservation between past and future boundary were shown to be equivalent to Low’s subleading soft photon theorem. While their expression for the current $\overset{2,0}{j_{A}}$ only includes free matter contributions, our expression accounts for infrared corrections from the coupling between the massless matter and the gauge field,
@@ 6 Conclusion
The long-range nature of electromagnetic interactions has been shown to modify the low-energy expansion of the S-matrix in scalar QED by logarithmic terms [Sahoo:2018lxl] which can be translated to a late-time divergence in the radiation dressing the matter by a logarithmically divergent phase and the universality of these results is explained by the underlying superphaserotation symmetry [Choi:2024mac]. The massless limit of the logarithmic soft photon theorem can immediately be seen to vanish (at least up to potential collinear divergences). This begs for an explanation in terms of symmetries.
@@ Appendix B Perturbatively corrected matter current
    $\overset{3,n}{j_{r}}$ $=ie\phi_{0,0}\big(\partial_{r}\overline{\phi}\big)_{1,n}+\text{c.c.}=-ie|\phi_{0,0}|^{2}\delta_{n,0}+\text{c.c.}=0.$ (B.5)
@@ Appendix D Absence of collinear divergences in the classical limit
The subscript ‘reg’ on the integral indicates that the integral is to be taken over the region where $|\ell^{\mu}|$ is large compared to $\omega$ but small compared to the energies of the finite energy particles, see [Sahoo:2018lxl] for details on the explicit evaluation of the integrals. Parameterizing the $\ell$-space in spherical coordinates, with $\vec{p}_{b}$ along the Z-axis and the XY-projection of $\vec{p}_{a}$ along the X-axis, calling the ‘fixed’ angle between $\vec{p}_{a}$ and $\vec{p}_{b}$ $\alpha_{ab}$, this integral becomes

### v2

@@ 1 Introduction
The simplest example of such an infrared triangle consists of the soft photon theorem in QED, the electromagnetic kick memory effect and large gauge symmetry which acts on matter as a local enhancement of global phase rotations, or, superphaserotations; it manifests itself in the conservation of a Noether charge at every angle whose global version is the familiar U(1) charge of QED [13, 20]. Analogous infrared triangles have been found in other theories including gravity [23], non-Abelian gauge theory [14] and supersymmetric theories [22]. The study of the infrared structure has also been extended to subleading orders in the low-energy expansion [20, 4, 15, 6] which translates to weaker large-distance fall-offs of the radiative field in the corresponding memory effect analyses and to overleading orders in a large-distance expansion of the associated symmetry parameters [5, 6, 9, 11].
The long-range nature of electromagnetic and gravitational interactions turns out to modify all these infrared triangles beyond the leading one: early/late-time divergences were shown to dress the matter fields [7] by logarithmically divergent phases and this behavior dominates over the free field asymptotics! Only recently have these effects been accounted for in the discussion of the infrared triangles [9, 10, 11] and so far only in (scalar) QED and gravity with massive matter. Our goal in this work is to connect the logarithmic soft photon theorem in massless scalar QED to an asymptotic symmetry analysis. A corresponding analysis for gravity coupled to massless matter is in progress [8].
@@
Unlike the leading soft factor (1.2), the subleading soft factor $S_{0}$ is not universal. It receives corrections from a (finite) set of effective field theory operators [12] with three-point interactions as well as loop corrections [2, 17, 18, 21]. These loop corrections enter at $O(\ln\omega)$ and so are overleading compared to the subleading tree-level soft factor $\sim\omega^{0}$ which the presence of the logarithm renders ambiguous.
@@
In (massive) scalar QED, it has been conjectured [16] that given a well-defined soft expansion to all orders in the loop expansion, the ratio between the $(N+1)$-particle amplitude including a soft photon and the corresponding $N$-particle amplitude without that soft photon is infrared-finite. This allows us to define the loop-corrected expansion of the soft factor as
@@
The logarithmic soft theorem in theories with massive matter receives contributions from classical infrared and quantum loop effects [21]
@@
An asymptotic symmetry explanation of the classical soft theorem was given in [10] which derived a ‘soft’ charge that maps on to the soft photon insertion in the S-matrix and a ‘hard’ charge whose action on the matter fields produces the classical soft factor. The conservation of the sum of these logarithmic soft and hard charges between past and future boundaries is exact in the electromagnetic coupling and yields the classical logarithmic soft photon theorem which is one-loop exact.
Here we would like to explain the symmetry underlying the classical logarithmic soft photon theorem in theories with massless matter. However, it is easy to see that the (naive [^2] The absence of potential collinear divergences is shown in appendix D.) massless limit of (1.7) vanishes. This points towards the classical one-loop corrected scattering structure of massless QED being trivial: long-range effects do not correct the subleading tree-level structure. Moreover the subleading tree-level soft theorem is not rendered ambiguous as they would have been in the presence of logarithms. The only non-vanishing contribution comes from quantum effects, but how to explain them from an asymptotic symmetry analysis is as of yet unclear, and we will not attempt to do so here [^3] Attempts in this direction were made in [7].. In an asymptotic symmetry derivation of the classical logarithmic soft theorem with massless matter we thus have to show that both soft and hard charges exactly vanish to all orders in the electromagnetic coupling!
@@
As ansatz for the relevant symmetry transformation we consider the same linearly divergent superphaserotation that is responsible for the subleading tree-level soft photon theorem. This is motivated by the corresponding derivation of the logarithmic soft photon theorem in massive scalar QED. Indeed we will see here that this ansatz reproduces the leading and subleading tree-level soft theorems in massless scalar QED in addition to explaining the exact vanishing of logarithmic soft photon theorem from symmetry [^4] See also the related work [1] .. Moreover, we determine the long-range correction to the charge associated to the subleading soft theorem [5], at higher order in the electromagnetic coupling which corresponds to the one-loop correction to Low’s subleading soft photon theorem. Care has to be taken with potential collinear divergences, which we show do, however, not affect our classical analysis.
@@ 2.2 Interacting fields
Because the correction to the Klein-Gordon equation is subleading compared to the free equation, the interacting scalar field is still dominated by the free behaviour, $\phi(x)=\phi^{\rm free}(x)+o(1/r)$, which implies $\phi_{0,n}=\delta_{n,0}\phi_{0,0}$ in (2.26). This behavior is in stark contrast to the situation in massive QED [10] where Coulombic modes give rise to a logarithmically divergent dressing for the massive scalar field. In massless QED, the matter dressing by Coulombic gauge modes only comes in at subleading order
@@ 3 Superphaserotation symmetry
are referred to as large if they give rise to non-vanishing Noether charges on the asymptotic phase space of the theory. To avoid any potential confusion with gauge redundancies, we follow [10] and refer to these physical transformations by how they act on matter fields, namely as a local enhancement of global phase rotations, or, superphaserotations.
@@
The solution to this equation that is relevant for our discussion is the one that diverges linearly in $r$ [5],
@@
where ${\overline{\epsilon}(\hat{x})}$ is a free function on $S^{2}$. Asymptotic charges with such divergent symmetry parameters have been found in [5] to account for Low’s subleading tree-level soft photon theorem [19]. Accounting for the long-range interactions, it has been shown [10] that the same symmetry transformation governs the logarithmic soft photon theorem in massive scalar QED. It is thus natural to assume that the same symmetry that governs the massless limit of Low’s soft theorem is also responsible for the logarithmic soft photon theorem in massless scalar QED.
@@ 4 Superphaserotation charge
To make the connection with soft theorems, it is useful to identify the ‘soft’ and ‘hard’ contributions to this charge (4.8), i.e. the contributions that will map onto the soft photon insertion in the S-matrix and the contributions coming from the superphaserotation of the matter fields which will map onto the soft factor. Interestingly, unlike in massive scalar QED, the term $\text{d}\epsilon\wedge\star{F}$ in (4.8) turns out to contribute not just to the soft charge but also to the hard charge. This is already the case for the charge constructed in [5], whose conservation was shown to be equivalent to the subleading tree-level soft photon theorem in massless scalar QED. Here we will find a similar result for the logarithmic soft photon theorem.
@@
As in the massive QED analysis [10] it turns out the tower of logarithmic terms in (4.11) vanishes. Thus, the charge (4.10), computed from the symplectic structure is linearly divergent in $r$ and $u$, logarithmically divergent in $r$, and potentially logarithmically divergent in $|u|$, as arising from the $1/|u|$ term in (4.11). To regulate the logarithmic divergences at large $r$ and large-$|u|$ we introduce a cutoff [^5] Note that choosing different cutoffs to regulate large-$r$ and large-$u$ divergences only affects the charge corresponding to the subleading tree-level soft theorem. This ambiguity is expected: logarithmic terms in the soft expansion render the power-law tree-level expansion (beyond the Weinberg pole) ambiguous. at $\ln\Lambda^{-1}$, similar to the case of massive QED [10]. In the following we will show that the three charges, $Q^{(-1)}$, $Q^{(0)}$ and $Q^{(\ln)}$, that we can extract from (4.10) are associated to, respectively, the leading Weinberg, subleading Low, and logarithmic Sahoo-Sen soft theorems.
@@ 4.3 Loop-corrected subleading charge $Q^{(0)}$
The covariant form of this expression matches the one obtained in [5] whose conservation between past and future boundary were shown to be equivalent to Low’s subleading soft photon theorem. While their expression for the current $\overset{2,0}{j_{A}}$ only includes free matter contributions, our expression accounts for infrared corrections from the coupling between the massless matter and the gauge field,
@@ 6 Conclusion
The long-range nature of electromagnetic interactions has been shown to modify the low-energy expansion of the S-matrix in scalar QED by logarithmic terms [21] which can be translated to a late-time divergence in the radiation dressing the matter by a logarithmically divergent phase and the universality of these results is explained by the underlying superphaserotation symmetry [10]. The massless limit of the logarithmic soft photon theorem can immediately be seen to vanish (at least up to potential collinear divergences). This begs for an explanation in terms of symmetries.
@@ Appendix B Perturbatively corrected matter current
    $\overset{3,n}{j_{r}}$ $=ie\phi_{0,0}\big(\partial_{r}\phi^{*}\big)_{1,n}+\text{c.c.}=-ie|\phi_{0,0}|^{2}\delta_{n,0}+\text{c.c.}=0.$ (B.5)
@@ Appendix D Absence of collinear divergences in the classical limit
The subscript ‘reg’ on the integral indicates that the integral is to be taken over the region where $|\ell^{\mu}|$ is large compared to $\omega$ but small compared to the energies of the finite energy particles, see [21] for details on the explicit evaluation of the integrals. Parameterizing the $\ell$-space in spherical coordinates, with $\vec{p}_{b}$ along the Z-axis and the XY-projection of $\vec{p}_{a}$ along the X-axis, calling the ‘fixed’ angle between $\vec{p}_{a}$ and $\vec{p}_{b}$ $\alpha_{ab}$, this integral becomes